OpenStudy (anonymous):

At a given time the height of a person is 5 feet and it increases at a rate of 1 inch per year. The person's weight is 100 lb and it increases at a rate of 5 lb/year. (a) What is the body fattness (BMI) of the person? (b) Is the person getting slimmer? (c) Assuming that the instantaneous rate of change of BMI does not change over a year, predict the person's BMI one year from now. Recall that the BMI index is the fraction of the body weight measured in kilograms divided by the square of the body height measure in meters.

OpenStudy (anonymous):

we have to convert weight in kg and height in meters and i did that , can u just check the answer ? a is 20.22 b is no and c is 20.33

OpenStudy (anonymous):

and isnt w= 100+5t

OpenStudy (anonymous):

Sorry, you are quite right, must convert and then re-write my equations,. 1 kg = 2.2 lb, 1 inch = 0.0254 m.

OpenStudy (anonymous):

dont be , can you please check the answers?

OpenStudy (anonymous):

BMI = [100 lb / 2.2 kg/lb)]/{60 inches x 0.0254 m/inch]^2 = 19.57 W = (100/2.2) + (5.0/2.2) t t in years W in kg H = (60 + 1 t)(.0254) H is in m Getting slimmer? Lower BMI? Evaluate BMI at t=1 and compare with BMI ay t=0. t=1 value will also answer part c.